Calculating the price elasticity of supply.

Beth is a college student who lives in San Francisco and provides math tutoring for extra cash. At a wage of $30 per hour, she is willing to tutor 6 hours per week. At $50 per hour, she is willing to tutor 16 hours per week.

Using the midpoint method, the elasticity of Beth's labor supply between the wages of $30 and $50 per hour is approximately ( ), which means that Beth's supply of labor over this wage range is ( ).
Answer #1

P1 = 30             P2 =50

Q1 = 6              Q2 = 16

Elasticity of supply = (16-6) / (50-30) * (50+30) / (6+16)

                            = (10/20) * (80/22) = 1.82

Using the midpoint method, the elasticity of Beth's labor supply between the wages of $30 and $50 per hour is approximately 1.82, which means that Beth's supply of labor over this wage range is elastic.